OpenStudy (anonymous):

Solve the indefinite integral cosx/ (sinx)^2 dx

OpenStudy (anonymous):

Let \(sin(x) = t\)

OpenStudy (anonymous):

\[\sin(x) = t\] \[\cos(x) dx = dt\] \[\int\limits \frac{dt}{t^2} = ??\]

OpenStudy (anonymous):

\[\int\limits_{}^{} \cos x / \sin ^{2}x \] I have that u=sinx, and du=cosx So then it looks like \[\int\limits_{}^{} du/u^2\]

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

-u^-1 + C ?

OpenStudy (anonymous):

\[\int\limits u^{-2} du \implies \frac{u^{-2 + 1}}{-2 + 1} = \frac{u^{-1}}{-1} + C\]

OpenStudy (anonymous):

Yes well done.. Now plug back for u ..

OpenStudy (anonymous):

-1/sinx +c

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Welcome dear..